OpenStudy (anonymous):

A line segment is drawn from (1, 6) to (3, 6) on a coordinate grid. Which answer explains one way that the length of this line segment can be determined? A. Add 3 + 1. B. Add 6 + 1. C. Subtract 3 – 1. D. Subtract 6 – 1.

OpenStudy (alexandervonhumboldt2):

C grath it to see

OpenStudy (alexandervonhumboldt2):

well i have a lot oiof expiirence so i dont need a whole grath i just imagine

OpenStudy (danjs):

well the points are the form (x,y) both y coordinates are the same, so this will be a horizontal line... can you see that?

OpenStudy (danjs):



OpenStudy (danjs):

So the distance between the two , is just a horizontal distance from x=1 to x=3 Distance AB = subtract 3 - 1

OpenStudy (danjs):

More formally , the distance magnitude is the hypotenuse of a right triangle \[\left| AB \right| \sqrt{(By-Ay)^2 + (Bx - Ax)^2}\] By = 6 Ay = 6 Ax = 3 Bx = 3 \[\left| AB \right| = \sqrt{(0 - 0)^2 + (3 - 1)^2} = \sqrt{4} = 2\]

OpenStudy (anonymous):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{green}{ \text{Welcome to Open Study! }}\color{red}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{Welcome to Open Study! }}\color{yellow}\bigstar}\)

OpenStudy (anonymous):

@Hannah456789

OpenStudy (anonymous):

Thanks